# ArchesClimate: Probabilistic Decadal Ensemble Generation With Flow Matching

### v1

@@ 3.5 CRPS and Variance of ArchesClimate
Table 2: Comparison of CRPS and Variance between ArchesClimate and IPSL-DCPP. Scores are averaged over space and time for three periods: 1969-1979, 1979-1989 and 2010-2020. Both metrics keep the units shown in the table.
@@ 3.6 Interchangeability using Rank Histograms
Ensemble members generated by ArchesClimate should be able to be exchanged with members of IPSL-DCPP. We take a 10-year 10-member IPSL-DCPP ensemble initialized at 1969 for the North Atlantic (0°–65° N, 80°–0° W) and compare a 10-member ensemble of ArchesClimate one at a time for the same region and period. See Section 3.1 for a more detailed explanation of rank histograms. We chose the North Atlantic to complement the earlier assessment of the Tropics, as the North Atlantic exhibits important dynamics at the decadal timescale that differ from the Tropics \@BBOPcitep\@BAP\@BBN(Eade et al., 2014)\@BBCP. We use the same variable subset as in Section 3.5.
ArchesClimate produces relatively flat rank histograms for net_flux, wap, hur and ta, which indicate that the rank of the ArchesClimate member is likely to appear at each rank of the IPSL-DCPP ensemble. Bias in a rank histogram is represented by a skew to either end of the rank histogram. ArchesClimate often under-predicted values in the ocean variables tos and thetaot2000 and therefore was often ranked as the lowest. These histograms are good evidence that for several variables, our ensemble members are exchangeable with the IPSL-DCPP ensemble.
@@ 3.7 Comparison of Temporal Power Spectra
This experiment compares the Temporal Power Spectra (TPS) of both IPSL-DCPP and ArchesClimate to compare signals at different temporal frequencies. We look at the TPS for a 10-year 10-member ensemble initialized at 1969 for the North Atlantic. We use anomalies for all variables and use the same variable subset as in Section 3.3.
@@
In Figure 6, there is a noticeable annual, seasonal and monthly signal for all variables. For all variables that have u-shaped rank histograms, ArchesClimate is underpowered in all but the strongest cycles. In psl (sea-level pressure), ArchesClimate correctly captures the strong annual cycle but under-represents everywhere else. This gives us insight into where ArchesClimate is unable to capture the variability of certain variables. It is clear in Figure 6 that the variables with lower variance are unable to generate enough power at frequencies outside of monthly, seasonal and annual cycles. This helps inform how to increase variance by including spectral information in the loss function. We investigate this idea in Section 3.10.
@@
Figure 6: Diagnostics for the North Atlantic for years 1969-1979. On the left, Rank Histograms for ArchesClimate showing normalized frequency of the rank of a single ArchesClimate member in a 10-member ensemble of IPSL-DCPP. On the right, comparison of the temporal power spectral density of anomalies across time for ArchesClimate and IPSL-DCPP. The x-axis is logarithmic, with the smallest value being one month.
3.8 Calculating Linear Trend with Pearson Correlation Coefficient
@@
We use Pearson Correlation Coefficient (PCC) to understand how well the ensemble mean of ArchesClimate captures spatial decadal trends of IPSL-DCPP. We calculate the per-pixel PCC for sea surface temperature anomalies, using a 10-year 5-member ensemble from ArchesClimate.
@@
We show in Figure 7 that in regions with strong teleconnections, i.e. the North Atlantic, Pacific, and Indian Ocean, similar correlation over the decade as IPSL-DCPP. The persistence of these correlations in key basins indicates that ArchesClimate retains a similar level of decadal-scale predictive skill as IPSL-DCPP. ArchesClimate decorrelates throughout the decade in the high latitudes much more than IPSL-DCPP. In each of the three test periods, there is a decorrelation near Northern Canada and Greenland. ArchesClimate does not represent any sea-ice or ice dynamics, and is therefore limited in its ability to capture long-term dynamics of Arctic regions.
@@
Figure 7: Pearson Correlation Coefficient (PCC) of tos (sea surface temperature) anomalies for several periods for both IPSL-DCPP and ArchesClimate. The PCC is calculated per pixel over the 10 years.
@@ 3.9 Spatial Anomaly Analysis in the North Atlantic
In previous experiments, we investigated the ability of ArchesClimate to capture regionally and globally averaged patterns (Section 3.3 and Section 3.5). We look again at the anomalies of sea surface temperature in the North Atlantic, but now to understand how well ArchesClimate captures the spatial signal of the seasonal cycle. We compute averages of seasonal sea surface temperature anomalies for the 10-year test period initialized at 1969 over the North Atlantic (0°–65° N, 80°–0° W). The seasonal means are MAM (March, April, May), JJA (June, July, August), SON (September, October, November) and DJF (December, January, February). We compare the results to the baseline outlined in Section 3.1.
Figure 8: Averages of seasonal anomalies of sea surface temperature over the North Atlantic (0°–65° N, 80°–0° W) for 1969-1979.
@@ 3.10 Improving Variance in ArchesClimate
In this experiment, we explore ways to increase the variance lacking in ArchesClimate shown in Table 2. We compare three methods to improve variability. First, we multiply the initial noise at inference time by 1.1 (referred to as AC_noised) following ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP. The generative model takes the scaled noise distribution and translates this to a higher variance output distribution. Second, we apply noise scaling proportional to the difference in variance between the generated data and the target data for the validation period for each variable (AC_per_variable). Finally, we train another model with a spectral and image gradient loss (AC_updated_loss). To balance the weighting of the loss functions, we scale the gradient loss and spectral loss by 0.2. The loss function is as follows:
@@
Figure 9: Variance (top) and CRPS (bottom) of different techniques to increase variance in ArchesClimate. AC_noised increases initial noise at inference time by a factor of $1.$ AC_updated_loss uses an alternate loss function including gradient and spectral components. AC_per_variable scales the initial noise by the difference of variance in a generated ensemble versus the target ensemble for the test period for each variable.
@@ 4 Discussion and Conclusions
While we train with data that is the output of a climate model, the model does not receive the initial conditions used to start each ensemble member of IPSL-DCPP. The impact of omitting the initialization state of the climate model is limited, as initialization noise has been shown to fade after a short period \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. In ArchesClimate, we create variation between ensemble members by generating states from different samples of Gaussian noise.
@@
We include a holistic climate state in order to improve forecasting accuracy and generate more physically consistent states. In future work, however, possible that ArchesClimate could benefit from a more thorough set of forcings and input variables (e.g. sea ice, land-use). Similarly, a running average of the last decade’s climate could help the model as a sort of memory, especially for variables that operate on longer timescales \@BBOPcitep\@BAP\@BBN(Mignot et al., 2016)\@BBCP. By better constraining the model through these extra variables, we would expect to produce a more constrained output.
@@
Our objective was to capture the underlying distribution of each physical variable at each timestep so that we can generate samples of the distribution of a given climate while still obeying dynamics that are resolved at shorter timescales. This task is difficult as the target distribution is hidden and the assessment of samples from the distribution requires expertise in climate science \@BBOPcitep\@BAP\@BBN(Mignot et al., 2016)\@BBCP. To aid the assessment of samples, we proposed a set of evaluations of both the statistics and the physical properties of the AI-generated ensembles that can be used in future evaluations of probabilistic climate model emulators. Further work can be done to find a base set of metrics and evaluations using already existing tools such as PCMDI, ESMValTool, climpred, and xMIP. With more thorough metrics, we can increase confidence in the ability of climate model emulators to augment climate models. Analyses can also be done to see if the climate model emulator adheres to conservation properties such as hydrostatic constraints and total water in atmospheric columns \@BBOPcitep\@BAP\@BBN(Sha et al., 2025; White et al., 2024; Watt-Meyer et al., 2024)\@BBCP.
@@
Our work opens several new directions of research. Using members generated from ArchesClimate, we believe it will be possible to temporally and spatially downscale using similar generative methods to enable analysis at higher resolutions. Further work also includes training on longer experiments such as the Coupled Model Inter-comparison Project experiment historical simulations that span several hundred years. By extending ArchesClimate to multi-decadal climate simulations, we can assess if it can emulate inter-annual variability at long timescales \@BBOPcitep\@BAP\@BBN(Jain et al., 2023)\@BBCP. Expanding the forcings already used in ArchesClimate, interpolation between longer experiments can be investigated to help climate scientists explore previously untested future climate scenarios.
@@
Another possibility is to explore recent advances in generative methods, for which inference time can be done in a fraction of the time, sometimes using only a single inference step \@BBOPcitep\@BAP\@BBN(Hess et al., 2025; Schmitt et al., 2024)\@BBCP, which would cut the cost of generating samples by an order of magnitude. Our results suggest that emulating climate states at a monthly resolution is an efficient way to predict long-term climate dynamics. Further work is needed to explore the limits of jointly predicting multiple dynamics at monthly time scales. It remains an open question whether ArchesClimate is primarily capturing the underlying processes that operate at shorter timescales, or whether some of its skill at monthly prediction may arise from correlations present at those scales.
@@ Appendix C Alternative Train/Test Splits
In the train/test split of IPSL-DCPP described in Section 2.3, there is temporal overlap in the training and test sets. Here, we look at a train/test split that leaves a test set temporally apart from the train set. To do this, we use the initialization years 1960-2000, so the last year of training is 2009, which we refer to in Figure 11 as alternative. We can then compare a model that has seen the years 2010-2020 to a model that has not. We generate a 10-member ensemble initialized in 2010 for both models.
@@
Figure 11: Comparison of CRPS for different train/test splits. “original” indicates the training scheme outlined in Section 2.3 and “alternative” indicates the training scheme where training and test contain no temporal overlap. The dotted line is the 5-member IPSL-DCPP baseline.
Open Research Section
Code for the project can be found at https://github.com/INRIA/geoarches. Training data can be found at https://esgf-node.ipsl.upmc.fr/projects/cmip6-ipsl/.
@@ Acknowledgements.
G. Clyne, G. Couairon, and C. Monteleoni were primarily supported by the Choose France Chair in AI, from the French government. The authors thank Juliette Mignot, David Landry, Clément Dauvilliers, and Renu Singh for several discussions about the project. To process the data from IPSL, this study benefited from the IPSL mesocenter ESPRI facility, which is supported by CNRS, UPMC, Labex L-IPSL, CNES and Ecole Polytechnique.

### v2

@@ 3.5 CRPS and Standard Deviation of ArchesClimate
Table 3: Comparison of CRPS and Standard Deviation between ArchesClimate, IPSL-DCPP and Pattern Scaling. Scores are averaged over space and time for three test periods: 1969-1979, 1979-1989 and 2010-2020. Both metrics keep the units shown in the table. All stats are generated from 5-member ensembles and are compared against a held-out 5-member ensemble from IPSL-DCPP. Closer to IPSL-DCPP is better.
@@ 3.6 Interchangeability using Rank Histograms
To evaluate whether ensemble members generated by ArchesClimate are exchangeable with those of IPSL-DCPP, we conducted a rank histogram analysis. We compared a 10-member ArchesClimate ensemble against a 10-member IPSL-DCPP ensemble, both initialized in 1969, over a 10-year period in the North Atlantic (0°–65° N, 80°–0° W). This region was selected to complement our earlier assessment of the Tropics, as the North Atlantic exhibits distinct decadal dynamics (Eade et al., 2014). We take a single member at a time from the ArchesClimate ensemble, and determine its pixel-wise rank among the 10-member IPSL-DCPP ensemble. We then average the rank of all ten ArchesClimate members. This gives us many samples to have a clear picture on the interchangeability of the ensemble members generated by ArchesClimate. We utilized the same variable subset as in subsection 3.5. See subsection 3.1 for a detailed methodology on rank histograms. The results indicate varying degrees of exchangeability. ArchesClimate produces relatively flat rank histograms for net_flux and wap, (see Figure 6), suggesting that ArchesClimate members are statistically indistinguishable from IPSL-DCPP members for these variables. However, we observed distinct biases in other areas. ArchesClimate consistently under-predicted values for tos and thetaot2000, resulting in a rank histogram skewed toward the lowest ranks. In the atmosphere, several variables exhibited n-shaped (convex) histograms, indicating that ArchesClimate members are more likely to rank in the middle of the IPSL-DCPP ensemble. This is consistent with the findings in subsection 3.5, which showed that ArchesClimate generally exhibits lower variance than IPSL-DCPP.
@@ 3.7 Comparison of Temporal Power Spectra
This experiment compares the Temporal Power Spectra (TPS) of both IPSL-DCPP and ArchesClimate to compare signals at different temporal frequencies. We look at the TPS of a 10-year 10-member ensemble initialized at 1969 for the North Atlantic, and take the mean across members. We use anomalies for all variables and use the same variable subset as in subsection 3.3.
@@
In Figure 6, there is a noticeable annual, seasonal and monthly signal for all variables in ArchesClimate. ArchesClimate is over-predicting the annual and seasonal cycle, suggesting that ArchesClimate is memorizing the annual and seasonal cycles. For all variables that have n-shaped rank histograms, ArchesClimate is underpowered in all but the strongest cycles. In psl (sea-level pressure), ArchesClimate correctly captures the strong annual cycle but under-represents everywhere else. This gives us insight into where ArchesClimate is unable to capture the variability of certain variables. It is clear in Figure 6 that the variables with lower variance are unable to generate enough power at frequencies outside of monthly, seasonal and annual cycles. This helps inform how to increase variance by including spectral information in the loss function. We investigate this idea in subsection 3.10.
@@
Figure 6: Diagnostics for the North Atlantic for test period 1969-1979. On the left, rank histograms for ArchesClimate showing normalized frequency of the rank of a each ArchesClimate member in a 10-member ensemble of IPSL-DCPP. Values closer to one indicate better interchangeability with IPSL-DCPP. On the right, comparison of the temporal power spectral density of anomalies of 10-member ensembles across time for ArchesClimate and IPSL-DCPP. The x-axis is logarithmic, with the smallest value being one month.
3.8 Calculating the Correlation of the Linear Trend
@@
We calculate the correlation of the linear trend to understand how well the ensemble mean of ArchesClimate captures spatial decadal trends of IPSL-DCPP. We calculate the per-pixel correlation of linear trend between different models for sea surface temperature anomalies. We use a 5-member ensemble for the periods 1969-1979,1979-1989 and 2010-2020 from ArchesClimate. For the same period, we use a 10-member ensemble from IPSL-DCPP, which we split the ensemble into two 5-member ensembles. First we use to compute the correlation of the two 5-member IPSL-DCPP ensembles (IPSL vs. IPSL in Figure 7). We then compute the correlation of the linear trend of the 5-member ArchesClimate ensemble and one of the 5-member IPSL-DCPP ensembles.
@@
We show in Figure 7 that in regions with strong teleconnections, i.e. the North Atlantic, Pacific, and Indian Ocean, ArchesClimate has similar linear trend over the decade as IPSL-DCPP. The persistence of these correlations in key basins indicates that ArchesClimate retains a similar level of decadal-scale predictive skill as IPSL-DCPP. ArchesClimate decorrelates throughout the decade in the high latitudes much more than IPSL-DCPP. In each of the three test periods, there is a decorrelation near Northern Canada and Greenland. ArchesClimate does not represent any sea-ice or ice dynamics, and is therefore limited in its ability to capture long-term dynamics of Arctic regions. The linear trend in 1979-1989 is less correlated than the other two decades, and is accordingly represented in ArchesClimate.
@@
Figure 7: Correlation of tos (sea surface temperature) anomalies for several periods for both IPSL-DCPP and ArchesClimate. We use the ensemble mean of 5-member ensembles. The correlation calculated by finding the linear trend of the anomalies between IPSL-DCPP and ArchesClimate per pixel over the 10 years. Values closer to one indicate higher correlation.
@@ 3.9 Spatial Anomaly Analysis in the North Atlantic
In previous experiments, we investigated the ability of ArchesClimate to capture regionally and globally averaged patterns (subsection 3.3 and subsection 3.5). We look again at the anomalies of sea surface temperature in the North Atlantic, but now to understand how well ArchesClimate captures the spatial signal of the seasonal cycle. We compute averages of seasonal sea surface temperature anomalies for the 10-year test period initialized at 1969 over the North Atlantic (0°–65° N, 80°–0° W). The seasonal means are MAM (March, April, May), JJA (June, July, August), SON (September, October, November) and DJF (December, January, February). We compare the results to the baseline outlined in subsection 3.1.
Figure 8: Averages of seasonal anomalies of sea surface temperature over the North Atlantic (0°–65° N, 80°–0° W) for the test period 1969-1979. Here we compare a 5-member ensemble from ArchesClimate to a 5-member ensemble of IPSL-DCPP seasonally averaged over the decade. This complements the statistics shown in subsection 3.5.
@@ 3.10 Improving Variance in ArchesClimate
In this experiment, we explore ways to increase the variance lacking in ArchesClimate shown in Table 3. We compare three methods to improve variability. First, we multiply the initial noise at inference time by 1.1 (referred to as AC_noised) following ArchesWeatherGen (Couairon et al., 2024). The generative model takes the scaled noise distribution and translates this to a higher variance output distribution. Second, we apply noise scaling proportional to the difference in variance between the generated data and the target data for the validation period for each variable (AC_per_variable). Finally, we train another model with a spectral and image gradient loss (AC_updated_loss). To balance the weighting of the loss functions, we scale the gradient loss and spectral loss by 0.2. The loss function is as follows:
@@
Figure 9: Variance (top) and CRPS (bottom) of different techniques to increase variance in ArchesClimate. AC_noised increases initial noise at inference time by a factor of $1.$ AC_updated_loss uses an alternate loss function including gradient and spectral components. AC_per_variable scales the initial noise by the difference of variance in a generated ensemble versus the target ensemble for the test period for each variable. We compare 5-member ensembles for the test period 1969-1979. Values closest to the IPSL-DCPP 5-member ensemble (the dotted black line) are best.
@@ 4 Discussion and Conclusions
While we train with data that is the output of a climate model, the model does not receive the initial conditions used to start each ensemble member of IPSL-DCPP. The impact of omitting the initialization state of the climate model is limited, as initialization noise has been shown to fade after a short period (Smith et al., 2019). In ArchesClimate, we create variation between ensemble members by generating states from different samples of Gaussian noise.
@@
We include a holistic climate state in order to improve forecasting accuracy and generate more physically consistent states. In future work, however, possible that ArchesClimate could benefit from a more thorough set of forcings and input variables (e.g. sea ice, land-use). Similarly, a running average of the last decade’s climate could help the model as a sort of memory, especially for variables that operate on longer timescales (Mignot et al., 2016). By better constraining the model through these extra variables, we would expect to produce a more constrained output.
@@
Our objective was to capture the underlying distribution of each physical variable at each timestep so that we can generate samples of the distribution of a given climate while still obeying dynamics that are resolved at shorter timescales. This task is difficult as the target distribution is hidden and the assessment of samples from the distribution requires expertise in climate science (Mignot et al., 2016). To aid the assessment of samples, we proposed a set of evaluations of both the statistics and the physical properties of the AI-generated ensembles that can be used in future evaluations of probabilistic climate model emulators. Further work can be done to find a base set of metrics and evaluations using already existing tools such as PCMDI, ESMValTool, climpred, and xMIP. With more thorough metrics, we can increase confidence in the ability of climate model emulators to augment climate models. Analyses can also be done to see if the climate model emulator adheres to conservation properties such as hydrostatic constraints and total water in atmospheric columns (Sha et al., 2025; White et al., 2024; Watt-Meyer et al., 2024).
Despite the encouraging results, this work is subject to several key limitations. Primarily, the use of coarse temporal resolution prevents fine-grained analysis of high-frequency climate phenomena. This modeling choice was motivated by the higher predictability of low-frequency phenomena, and the computational gains of predicting climate states at a monthly temporal resolution. Furthermore, ArchesClimate directly inherits any inherent biases present in the underlying climate model, IPSL-CM6A-LR. A significant technical constraint observed is the underprediction of variance across several critical variables, which may limit the model’s ability to capture extreme events. Additionally, the variable set is limited and contains a restricted representation of ocean variables, thereby constraining the scope of oceanic climate interactions that can be studied. We trained a single model on a short time period that contained a large amount of data to better evaluate the model’s learning limits. Finally, the ability to conduct comprehensive model ablations and hyperparameter sensitivity analyses was constrained by the high computational resource requirements necessary for training and evaluation.
@@
Our work opens several new directions of research. Using members generated from ArchesClimate, we believe it will be possible to temporally and spatially downscale using similar generative methods to enable analysis at higher resolutions. Further work also includes training on longer experiments such as the Coupled Model Inter-comparison Project experiment historical simulations that span several hundred years. By extending ArchesClimate to multi-decadal climate simulations, we can assess if it can emulate inter-annual variability at long timescales (Jain et al., 2023). Expanding the forcings already used in ArchesClimate, interpolation between longer experiments can be investigated to help climate scientists explore previously untested future climate scenarios.
@@
Another possibility is to explore recent advances in generative methods, for which inference time can be done in a fraction of the time, sometimes using only a single inference step (Hess et al., 2025; Schmitt et al., 2024), which would cut the cost of generating samples by an order of magnitude. Our short investigation into the ability of ArchesClimate to stably generate states for longer temporal periods (see Appendix B suggest that emulating climate states at a monthly resolution is an efficient way to predict long-term climate dynamics. Further work is needed to explore the limits of jointly predicting multiple dynamics at monthly time scales. It remains an open question whether ArchesClimate is primarily capturing the underlying processes that operate at shorter timescales, or whether some of its skill at monthly prediction may arise from correlations present at those scales.
@@ Appendix C Alternative Train/Test Splits
In the train/test split of IPSL-DCPP described in subsection 2.4, there is temporal overlap in the training and test sets. Here, we look at a train/test split that leaves a test set temporally apart from the train set. To do this, we use the initialization years 1960-2000, so the last year of training is 2009, which we refer to in Figure 11 as alternative. We can then compare a model that has seen the years 2010-2020 to a model that has not. We generate a 10-member ensemble initialized in 2010 for both models.
@@
Figure 11: Comparison of CRPS for different train/test splits. “original” indicates the training scheme outlined in subsection 2.4 and “alternative” indicates the training scheme where training and test contain no temporal overlap. Better values are closer to the dotted line, which is the 5-member IPSL-DCPP baseline. We compare 5-member ensembles for the test period 1969-1979.
@@ Acknowledgements.
[…]